The Model

quasistatic elastic 3D geometry fault patch (small dislocations with an associated area) Green’s function stress transfer event sequences

The Interface

constructing and editing model geometry and physical parameters running simulations, saving/reloading simulation state graphically displaying observables (stress, slip, failures)

Larger Issues

modularization modification development versus long runs web-based interaction standardization versus flexibility interfaces to programming environments